OpenStudy (anonymous):

one gallon equals 4 quarts. write a rule that expresses the number of quarts (q) as a function of the number of gallons (g)

OpenStudy (bahrom7893):

1g=4q q=(1/4)g

OpenStudy (anonymous):

now i need to : use your function from part a to find the number of quarts in 17 gallons

OpenStudy (bahrom7893):

q=(1/4)*17=17/4=4.25quarts

OpenStudy (anonymous):

thanks soo much